A certain color emerges at 15° in the first-order spectrum. At what angle would this same color
emerge in the second order if the same source and grating are used?
1
Expert's answer
2020-10-13T07:12:57-0400

Solution

Using following equation

"n\lambda=a sin\theta" ................. eq. 1

For first case using

n=1, "\theta=15^°"

"\frac{\lambda}{a}=sin 15^°=0.2588"


Now for second all things are same "\lambda" and a, n=2 and "\theta" is finding through equation 1

"2\lambda=asin\theta"

"\sin\theta=\frac{2\lambda}{a}=2\times0.2588"

"\theta=" "\arcsin(0.518) =31.17^°"

Therefore angle is "31.17^\u00b0"